# Enhance-A-Video: Better Generated Video for Free

### v2

@@ 1 Introduction
Diffusion transformer (DiT) models (Peebles & Xie, 2022) have revolutionized video generation, enabling the creation of realistic and compelling videos (Yang et al., 2024; Brooks et al., 2024; Lin et al., 2024; Kong et al., 2025). However, achieving temporal consistency across frames while maintaining fine-grained details remains a significant challenge. Many existing methods generate videos that suffer from unnatural transitions and degraded quality as illustrated in Figure 2, which fundamentally limits their practical applicability in real-world scenarios and professional applications (Yan et al., 2023; Henschel et al., 2024).
@@
Video generation enhancement (He et al., 2024) is designed for addressing the above limitations, where two objectives are primarily considered: ($i$) maintaining temporal consistency across frames, which ensures smooth and coherent transitions, and ($ii$) improving spatial details, which enhances the visual quality of each frame. In UNet-based video generation (Zhang et al., 2023a; Guo et al., 2024; Xu et al., 2024; Li et al., 2025), Upscale-A-Video (Zhou et al., 2024) integrated a local-global temporal strategy for better temporal coherence, and VEnhancer (He et al., 2024) designed a video ControlNet (Zhang et al., 2023b) to enhance spatial and temporal resolution simultaneously. Nevertheless, the exploration of enhancing DiT-based video generation remains limited, particularly in addressing challenges of temporal consistency and spatial detail preservation.
@@ 3.3 Temperature in DiT-based Video Generation
In video generation, a similar temperature principle can be considered when using DiT models, where the temporal attention mechanism controls the relationship between generated frames. A lower temperature enforces stronger frame-to-frame consistency by increasing temporal attention near the diagonal. In this case, the content structure exhibits high local temporal coherence, where each frame is correlated with its immediate neighbors and self-information. However, the extremely low temperature can bias the model toward the current frame, potentially suppressing diversity and causing repeated transitions in dynamic scenes.
Conversely, higher temperatures amplify non-diagonal temporal attention, allowing the DiT model to draw global information from multiple frames during generation. While this can enhance temporal diversity, setting the temperature extremely high results in uniform attention across all frames, possibly generating unexpected content or disrupting temporal coherence of the video output.
Equation 5 presents a direct usage of temperature in DiT. However, video generation requires a proper balance between cross-frame and intra-frame attention, if we directly apply the LLM-style temperature adjustment similarly to change attention weights , we always fail to enhance target cross-frame dependencies suitably.
@@
Directly applying $\tau$ to temporal attention causes increasing changes as the model deepens and denoising steps accumulate, which can lead to overly smooth motion, loss of visual details, and unstable video generation, as illustrated in Appendix A.
@@ 3.4 Enhance Block
$\textit{CFI}_{\textit{enhanced}}=\textbf{clip}(\tau\cdot\textit{CFI},1).$ (7)
@@ Appendix A Temperature Method Comparison
Figure 13: Temporal attention difference maps and corresponding generated videos comparing three temperature enhancement methods. (a) Temperature Attention Scaling $\tau=1$. (b) CFI Attention Scaling. (c) Enhance-A-Video Method.

### v3

@@ 1 Introduction
Diffusion transformer (DiT) models (Peebles & Xie, 2022) have revolutionized video generation, enabling the creation of realistic and compelling videos (Yang et al., 2024; Brooks et al., 2024; Lin et al., 2024; Xu et al., 2024a; Kong et al., 2025). However, achieving temporal consistency across frames while maintaining fine-grained details remains a significant challenge. Many existing methods generate videos that suffer from unnatural transitions and degraded quality as illustrated in Figure 2, which fundamentally limits their practical applicability in real-world scenarios and professional applications (Yan et al., 2023; Henschel et al., 2024).
@@
Video generation enhancement (He et al., 2024) is designed for addressing the above limitations, where two objectives are primarily considered: ($i$) maintaining temporal consistency across frames, which ensures smooth and coherent transitions, and ($ii$) improving spatial details, which enhances the visual quality of each frame. In UNet-based video generation (Zhang et al., 2023a; Guo et al., 2024; Xu et al., 2024b; Li et al., 2025), Upscale-A-Video (Zhou et al., 2024) integrated a local-global temporal strategy for better temporal coherence, and VEnhancer (He et al., 2024) designed a video ControlNet (Zhang et al., 2023b) to enhance spatial and temporal resolution simultaneously. Nevertheless, the exploration of enhancing DiT-based video generation remains limited, particularly in addressing challenges of temporal consistency and spatial detail preservation.
@@ 3.3 Temperature in DiT-based Video Generation
In video generation, a similar temperature principle can be considered when using DiT models, where the temporal attention mechanism controls the relationship between generated frames. Equation 5 presents a direct usage of temperature in temporal attention of DiT models.
@@
In particular, properly increased temperatures amplify non-diagonal temporal attention, allowing the DiT model to draw global information from multiple frames during generation, leading to better spatial diversity and temporal consistency. On the other hand, setting the temperature extremely high results in uniform attention across all frames, possibly generating unexpected or prompt-irrelevant content.
However, video generation requires a proper balance between cross-frame and intra-frame attention, if we directly apply the LLM-style temperature adjustment similarly to change the original attention weights , we always fail to enhance target cross-frame dependencies suitably. Directly applying $\tau$ to temporal attention causes increasing changes as the model deepens and denoising steps accumulate, which can lead to overly smooth motion, loss of visual details, and unstable video generation, as illustrated in Appendix A.
@@ 3.4 Enhance Block
$\textit{CFI}_{\textit{enhanced}}=\textbf{clip}((\tau+F)\cdot\textit{CFI},1).$ (7)
@@ Appendix A Temperature Method Comparison
Figure 13: Temporal attention difference maps and corresponding generated videos comparing three temperature enhancement methods. (a) Temperature Attention Scaling $\tau=1.1$. (b) CFI Attention Scaling. (c) Enhance-A-Video Method.